The average of 4 children is 10.the ages of 3 are 10,12,8.What is the age of the 4th child?
Answer by stanbon(75887)

The average of 4 children is 10.the ages of 3 are 10,12,8.What is the age of the 4th child?
----
average = sum of numbers/number of numbers
-----
Equation:
(x + 10 + 12 + 8)/4 = 10
----
x + 30 = 40
x = 10 (age of the 4th child)
===================================
Cheers,
Stan H.
==============